# Intelligence at the Edge of Chaos

### v2

@@ Abstract
We explore the emergence of intelligent behavior in artificial systems by investigating how the complexity of rule-based systems influences the capabilities of models trained to predict these rules. Our study focuses on elementary cellular automata (ECA), simple yet powerful one-dimensional systems that generate behaviors ranging from trivial to highly complex. By training distinct Large Language Models (LLMs) on different ECAs, we evaluated the relationship between the complexity of the rules’ behavior and the intelligence exhibited by the LLMs, as reflected in their performance on downstream tasks. Our findings reveal that rules with higher complexity lead to models exhibiting greater intelligence, as demonstrated by their performance on reasoning and chess move prediction tasks. Both uniform and periodic systems, and often also highly chaotic systems, resulted in poorer downstream performance, highlighting a sweet spot of complexity conducive to intelligence. We conjecture that intelligence arises from the ability to predict complexity and that creating intelligence may require only exposure to complexity.
@@ 1 Introduction
To investigate this hypothesis, we utilize Stephen Wolfram’s elementary cellular automata (ECA) as our experimental framework. ECAs are one-dimensional, binary-state, discrete computational systems defined by 256 possible 8-bit rules. They generate a diverse spectrum of behaviors ranging from simple, repetitive patterns to highly complex and chaotic structures (Wolfram, 1983). Despite their simple rule-based definitions, certain ECAs produce patterns of significant complexity, making them ideal for examining the relationship between intelligence and complexity.
Our methodology involves training separate instances of the GPT-2 language model (Radford et al., 2019) on datasets generated by individual ECAs. The models are tasked with predicting future states of the automata. Following this pretraining phase, we evaluate the models’ “intelligence” by quantifying performance on downstream logical reasoning and chess move prediction tasks.
This paper presents an extensive study exploring the relationship between system complexity and the emergence of intelligence in large language models (LLMs). We discover a positive correlation between the complexity of the ECA rules and the downstream performance of models trained on them, highlighting the role of complexity in fostering intelligent behavior. Surprisingly, we find that models can learn complex solutions when trained on simple rules. Our results point to an optimal complexity level, or “edge of chaos”, conducive to intelligence, where the system is structured yet challenging to predict. These findings enhance our understanding of intelligence in artificial systems and provide a framework for future research focused on the importance of complexity in developing these systems.
@@ 2.1 Elementary Cellular Automata
Cellular automata (CAs) are computational models of complex systems, consisting of a grid of cells that evolve over time based on simple rules. They were first introduced by John von Neumann in the 1940s (von Neumann, 1966). CAs have been widely used to simulate various physical, biological, and computational systems due to their simplicity and ability to produce complex behavior.
@@ 2.2 Large Language Models
Incorporating synthetic data into the training of large language models (LLMs) has been shown to significantly boost their performance. Synthetic data generation enables the creation of large, diverse, and high-quality datasets that may not be readily available in the real world (Anaby-Tavor et al., 2019). By augmenting training data with synthetic examples, models can learn from a broader range of linguistic patterns and rare occurrences, improving their ability to generalize to unseen scenarios (Edunov, 2018). As a result, LLMs trained with synthetic data exhibit enhanced robustness, accuracy, and adaptability across various downstream tasks (Feng et al., 2021).
@@ 2.3 Complexity Measures
While these measures are correlated with one another, each measures different aspects of complexity. For most analyses, we focus on Lempel-Ziv Complexity and the Wolfram Classification. Additional complexity measures are shown for performance on downstream tasks (Section  5.1).
@@ 3 Methodology
In this study, we systematically investigate the relationship between system complexity and emergent intelligence. This section outlines our methodology, including the steps for data generation and model pretraining. An overview of the training process and task evaluations is provided in Figure 1.
Figure 1: Our framework for investigating the link between complexity and intelligence. We pretrain Large Language Models (LLMs) on Elementary Cellular Automata (ECAs) from different complexity classes using next-token prediction, then evaluate them on downstream reasoning and chess move prediction tasks. We use various measures to analyze the complexity of ECA rules, and quantify the relationship between complexity and downstream performance.
@@ 3.1 Data Generation
To train our models, we simulate a selection of ECA rules. Each simulation generates a sequence of binary vectors, where each vector represents the system’s state at a specific time step. For each sample, we begin with a randomly initialized vector as the automaton’s initial state. The system is then evolved over 1000 time steps by repeatedly applying the chosen ECA rule. This process produces a sequence of binary vectors that capture the evolving dynamics of the ECA over time.
To increase the diversity of the training data, we extract random spatiotemporal windows from the full sequences. Specifically, we sample subsequences by selecting random windows of 60 time steps and 100 spatial dimensions from the binary vectors. This method exposes the model to a variety of contexts and state configurations, enhancing its ability to learn the dynamics of the ECA rules and generalize to new sequences. Each training sequence represents a randomly selected segment in space and time of the automaton’s evolution. We train models predicting either 1 or 5 steps in the future, to vary the difficulty of the task.
@@ 3.2 Training Procedure for GPT-2 Models
We utilized a modified GPT-2 architecture (Radford et al., 2019) adapted for binary input and output data, enabling it to perform next-token prediction on sequences of binary states. Instead of using a traditional token embedding layer followed by a softmax over a vocabulary, we replaced the token embeddings with a linear projection layer that directly maps binary vectors into the model’s embedding space. The GPT-2 model processes these embeddings to capture temporal dependencies and patterns within the sequences. At the output, we apply a linear projection layer to map the model’s hidden states back to the data dimensionality, generating the prediction for the next state of binary variables at each time step. This adaptation allows the GPT-2 model to handle binary data directly and perform next-token prediction without relying on a predefined vocabulary. This also makes the model deterministic, in line with the deterministic nature of ECAs.
@@ 3.3 Pretraining Setup
Each model was pretrained on next-token prediction using data generated from a single ECA rule for up to 10,000 epochs. Early stopping based on validation loss was utilized to prevent overfitting and conserve computational resources. The training data were organized into batches of 64 sequences, each comprising 60 time steps and 100 spatial dimensions.
@@ 4 Experiments
To evaluate the emergent intelligence of models trained on cellular automata, we conducted experiments on three downstream tasks: one easy and one hard reasoning task inspired by the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019), and a challenging chess move prediction task (Ruoss et al., 2024). These tasks were designed to quantify the models’ abilities in reasoning, abstraction, and long-term prediction, thereby assessing the level of intelligence encoded during pretraining on ECA rules of varying complexities. We freeze the layers of the pretrained GPT-2 models and train only the input and output projection layers for downstream tasks to ensure that performance differences reflect the inherent capabilities of the models.
Our central hypothesis is that models pretrained on complex rules will exhibit superior performance on downstream tasks compared to those pretrained on simple rules. The inclusion of both easy and hard tasks allows us to observe different performance trends and better understand the relationship between pretraining complexity, task difficulty, and emergent intelligence. The chess move prediction task, in particular, serves as an excellent system to test reasoning due to its inherent complexity and requirement for strategic thinking.
@@ 4.1 Downstream Task: Reasoning
Easy
In the easy task, we only use $3\times 3$ squares that are fixed in position and orientation. The only possible transformation is a color change, which we perform in a predetermined order.
Hard
The hard task involves more complex patterns where all transformations are applied simultaneously. We used four distinct base shapes, each represented by a $5\times 5$ matrix. This results in complex sequences where shapes change color, orientation, and position over time. The combination of these transformations requires the model to reason over multiple simultaneous changes to accurately predict the next pattern in the sequence.
Data Preparation and Training on the Downstream Tasks
For the easy reasoning task, models were trained for 1,000 epochs with a learning rate of $1\times 10^{-4}$. For the hard reasoning task, due to its increased complexity and difficulty, we post trained for 10,000 epochs with a learning rate of $1\times 10^{-5}$. Early stopping based on validation loss was again used to prevent overfitting.
@@ 4.2 Downstream Task: Chess Move Prediction
We added an embedding layer to convert the SAN tokens into vector representations, which were then processed using the frozen ECA-pretrained model. A linear output layer was used to transform the outputs to the vocabulary size corresponding to the SAN tokens. These input and output layers were trained while the rest of the model was frozen. The model was trained using cross-entropy loss, the Adam optimizer (Kingma, 2014), and a learning rate scheduler with warm-up. Early stopping was also implemented.
@@ 4.3 Hardware and Software
Figure 2: Relationship between downstream task performance and rule complexity. Left: Eight representative ECA rules, two from each of Wolfram’s four complexity classes. Performance of models trained on these rules is highlighted in the top row. Top: Model performance in relation to the Lempel-Ziv complexity of each rule. The left and center panels show efficiency (1 divided by number of epochs to reach 80% validation accuracy) for the easy and hard reasoning tasks, respectively. The right panel shows move prediction accuracy for the chess task. The rules depicted on the left are highlighted in the plot with triangles and annotated with the rule number. The correlation coefficient is shown in the top-left corner of each plot. An asterisk next to the value indicates a significant relationship ($p<0.05$). Bottom: Downstream task performance based on Wolfram classification of each rule. Models trained on chaotic and complex rules perform better than models trained on uniform and simple rules.
@@ 5 Results
In this section, we present our results exploring the relationship between system complexity and emergent intelligence in LLMs. The following sections detail our analyses of task performance and attention patterns across models trained on varying rule complexities.
@@ 5.1 Relationship between Intelligence and Complexity
Figure 3: Relationship between downstream task performance and rule complexity for other complexity measures. Rows depict easy reasoning, hard reasoning, and chess move prediction tasks, while columns show compression complexity, Lyapunov exponent, and Krylov complexity, respectively. We observe the same general patterns that we see with the Lempel-Ziv complexity in Figure 2.
Figure 2 presents the model performance across three downstream tasks (easy reasoning, hard reasoning, and chess move prediction) as a function of the complexity of the ECA rules the models were pretrained on. The top row highlights the relationship between performance and the Lempel-Ziv complexity of the rules, while the bottom row categorizes the performance by Wolfram’s complexity classes. For clarity, two representative rules from each complexity class are displayed on the left, with their corresponding performance annotated in the top plots.
For the reasoning tasks, models generally achieve near-perfect accuracy when trained for sufficient long time. Therefore, instead of reporting absolute accuracy, we focus on model efficiency, defined as the inverse of the number of epochs required to reach 80% accuracy. The chess task is sufficiently difficult that models do not achieve perfect performance, and so we report the final accuracy. As rule complexity increases, we observe a clear positive correlation in all tasks, with more complex rules leading to greater efficiency. This correlation is significant for each of the tasks ($p<0.05$).
In terms of Wolfram’s classification, rules from Classes I and II (uniform and periodic) show lower average efficiency in the reasoning tasks compared to those from Classes III and IV (chaotic and complex). Complex rules especially outperform the other classes on the chess move prediction task. This pattern suggests that models trained on more complex rules tend to perform better on harder downstream tasks. Results with respect to other complexity measures are shown in Figure 3.
We observe that models trained on certain Class III (Chaotic) rules, such as Rules 105, 146, and 150, have poorer performance on the hard reasoning and chess move prediction tasks. This behavior is expected due to chaotic systems lacking the structured patterns necessary for effective learning. In other words, they may be too random to predict, leading to weaker downstream performance. These results highlight the existence of a “sweet spot” of complexity conducive to intelligence, where the system is still predictable yet hard to predict.
@@ 5.2 Models Learn Complex Solutions For Simple Rules
Figure 4: Attention scores for the final 10 states prior to the target state, showing that models trained on more complex rules rely more heavily on past states for prediction. Left: Visualization of the last 10 states and the target state for representative rules from each of Wolfram’s complexity classes. Center: Attention scores for each of the last 10 states, highlighting that models trained on complex and chaotic rules focus more on recent states, while models trained on uniform rules exhibit consistently low attention. Periodic rules demonstrate a repeating attention pattern, suggesting that the model is learning to attend to earlier cycles of the same state rather than general state history. Right: Average attention across the final 10 states for all rules, plotted against Lempel-Ziv complexity. A strong positive correlation (r=0.66) indicates that models trained on higher complexity rules attend more highly to historical states during prediction.
@@
Our findings reveal that models trained on rules that produce more complex dynamics tend to allocate higher attention to the last ten states, with a strong positive correlation (r=0.66) between rule complexity and average attention. This suggests that models trained on complex rules integrate information from past states to make their predictions. In contrast, models trained on simpler rules display lower attention across the last ten states. For example, Rule 168 (uniform) shows consistently low attention, indicating that previous states are not being utilized in the prediction process. Rule 179 (periodic) demonstrates a recurring pattern in its attention scores, where the model focuses on every other state. This behavior is expected by the nature of Rule 179, which produces an alternating pattern that repeats every two time steps (Figure 4). Thus, the model appears to be learning only this alternating cycle, rather than general state history. The simpler attention structure suggests that these models learn a trivial solution, which is in line with their downstream performance.
@@
We emphasize that for both simple and complex rules, a trivial solution exists: learning the 8-bit instantaneous rule and applying it to only the current state. Such a model would have no attention on previous states, as it only needs the current state to make its prediction. The fact that the complex models are attending to previous states indicate that they are learning a more complex solution to this simple problem, and we conjecture that this complexity is what makes the model “intelligent” and capable of repurposing learned reasoning to downstream tasks.
@@
Figure 5: Performance comparison between models trained on short-term and long-term prediction tasks. Each point represents an ECA rule, with the x-axis showing the performance of models trained on short-term (1-step) prediction and the y-axis showing performance of models trained on long-term (5-step) prediction. The points are colored by the Lempel-Ziv complexity of each rule. The dashed line represents equal performance between the two models. Points below the line indicate better performance for short-term prediction models.
@@
We had initially expected that predicting one step in the future would be too easy, and would result in every model learning the trivial (8-bit rule) solution. As such, we trained models to predict 5 steps ahead. Surprisingly, we found that models predicting only the next step not only learned non-trivial solutions, but even outperformed models predicting five steps ahead (Figure 5). This result suggests that even when predicting the immediate next state, the models are learning nontrivial solutions (provided that the underlying rule is complex), capable of capturing complex patterns beyond the trivial rule-based approach.
@@
In this section, we explored the relationship between system complexity and emergent intelligence. Our results show that downstream model performance improves with pretraining on more complex rules, but can deteriorate with excessive complexity or chaotic behavior, signifying a sweet spot of complexity conducive to intelligence. Attention patterns further reveal that models trained on complex rules integrate historical information into their predictions, suggesting that they are learning more sophisticated solutions to relatively simple problems. We hypothesize that this complexity in the learned representations is a key factor enabling models to generalize and perform well on downstream tasks.

### v3

@@ Abstract
We explore the emergence of intelligent behavior in artificial systems by investigating how the complexity of rule-based systems influences the capabilities of models trained to predict these rules. Our study focuses on elementary cellular automata (ECA), simple yet powerful one-dimensional systems that generate behaviors ranging from trivial to highly complex. By training distinct Large Language Models (LLMs) on different ECAs, we evaluated the relationship between the complexity of the data generated by the rules and the models’ ability to learn effective general representations, as reflected in their performance on downstream tasks. Our findings reveal that models trained on more complex data exhibit greater predictive ability, as demonstrated by their performance on reasoning and chess move prediction tasks. Both uniform and periodic systems, and often also highly chaotic systems, resulted in poorer downstream performance, highlighting a sweet spot of complexity conducive to intelligence. We conjecture that intelligence arises from the ability to predict complexity and that creating intelligence may require only exposure to complexity.
11footnotetext: Equal contribution22footnotetext: Correspondence to david.vandijk@yale.edu
@@ 1 Introduction
To investigate this hypothesis, we utilize Stephen Wolfram’s elementary cellular automata (ECA) as our experimental framework. ECAs are one-dimensional, binary-state, discrete computational systems defined by 256 possible 8-bit rules. They generate a diverse spectrum of behaviors ranging from simple, repetitive patterns to highly complex and chaotic structures (Wolfram, 1983). Despite their simple rule-based definitions, certain ECAs produce significantly complex patterns, making them ideal for examining the relationship between complexity and intelligence.
Our methodology involves training separate instances of the GPT-2 language model  (Radford et al., 2019) on data generated by individual ECAs. The models are tasked with predicting future states of the automata. Following this pretraining phase, we evaluate the transformer models’ ability to learn useful representations by quantifying their performance on downstream logical reasoning and chess move prediction tasks.
This paper presents an extensive study exploring the relationship between system complexity and the emergence of intelligence in large language models (LLMs), quantified by the effectiveness of learned representations for downstream tasks. We discover a positive correlation between the complexity of the ECA data and the downstream performance of models trained on that data, highlighting the role of complexity in learning effective representations. Surprisingly, we find that models can learn complex solutions even when trained on rules that generate simple data. Our results suggest an optimal complexity level, or “edge of chaos”, conducive to learning, where the system is structured yet challenging to predict. These findings enhance our understanding of intelligence in artificial systems and provide a framework for future research focused on the importance of complexity in developing these systems.
@@ 2.1 Elementary Cellular Automata
Cellular automata (CAs) are computational models of complex systems, consisting of a grid of cells that evolve over time based on simple rules. First introduced by John von Neumann (von Neumann, 1966), CAs have since been widely used to simulate various physical, biological, and computational systems due to their simplicity and ability to produce complex behavior.
@@ 2.3 Complexity Measures
While these measures are correlated with one another, each measures different aspects of complexity. For most analyses, we focus on Lempel-Ziv Complexity and the Wolfram Classification. Performance on downstream tasks as a function of other complexity measures is shown in Appendix E.
@@ 3 Methodology
In this study, we systematically investigate the relationship between system complexity and the generalization of learned representations of models trained on these systems. This section outlines our methodology, including the steps for data generation and model pretraining. An overview of the training process and task evaluations is provided in Figure 1.
Figure 1: Our framework for investigating the link between complexity and intelligence. We pretrain Large Language Models (LLMs) on Elementary Cellular Automata (ECAs) from different complexity classes using next-token prediction, then evaluate them on downstream reasoning and chess move prediction tasks. We use various measures to analyze the complexity of ECA-generated data, and quantify the relationship between complexity and downstream performance.
@@ 3.1 Data Generation
To train our models, we simulate a selection of ECA rules. Each simulation generates a sequence of binary vectors, where each vector represents the system’s spatial state at a specific time step. For each sample, we begin with a randomly initialized vector as the automaton’s initial state. The system is then evolved over 1000 time steps by repeatedly applying the chosen ECA rule. This process produces a sequence of binary vectors that capture the evolving dynamics of the ECA over time.
To increase the diversity of the training data, we extract random spatiotemporal windows from the full sequences. Specifically, we sample subsequences by selecting random windows of 60 time steps and 100 spatial dimensions from the binary vectors. This method exposes the model to a variety of contexts and state configurations, enhancing its ability to learn the dynamics of the ECA rules and generalize to new sequences. Each training sequence represents a randomly selected segment in space and time of the automaton’s evolution. To vary the difficulty of the task, we train models to predict either 1 or 5 steps in the future.
@@ 3.2 Training Procedure for GPT-2 Models
We utilized a modified GPT-2 architecture (Radford et al., 2019) adapted for binary input and output data, enabling it to perform next-token prediction on sequences of binary vectors. Instead of using a traditional token embedding layer followed by a softmax over a vocabulary, we replaced the token embeddings with a linear projection layer that directly maps binary vectors into the model’s embedding space. The GPT-2 model processes these embeddings to capture temporal dependencies and patterns within the sequences. At the output, we apply a linear projection layer to map the model’s hidden states back to the data dimensionality, generating the prediction for the next state of binary variables at each time step. This adaptation allows the GPT-2 model to handle binary data directly and perform next-token prediction without relying on a predefined vocabulary. Additionally, this also makes the model deterministic, in line with the deterministic nature of ECAs.
@@ 3.3 Pretraining Setup
Each model was pretrained on next-token prediction tasks using data generated from a single ECA rule for up to 10,000 epochs. In each epoch, a new dataset was generated from a random initial state. Not only does this setup effectively provide infinite data, but it also makes the volume of data seen by the model directly proportional to the number of epochs the model has been trained for. To prevent overfitting, early stopping based on validation loss was employed. The training data were organized into batches of 64 sequences, each comprising 60 time steps and 100 spatial dimensions.
@@ 4 Experiments
To evaluate the emergent intelligence of models trained on cellular automata, we conducted experiments on three downstream tasks: one easy and one hard reasoning task inspired by the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019), and a challenging chess move prediction task (Ruoss et al., 2024). These tasks were designed to quantify the transformer models’ abilities in reasoning, abstraction, and long-term prediction, thereby assessing the level of intelligence encoded during pretraining on ECA-generated data of varying complexities. We freeze the layers of the pretrained GPT-2 models and train only the input and output projection layers for downstream tasks to ensure that performance differences reflect the inherent capabilities of the models.
Our central hypothesis is that models pretrained on rules that generate complex data will exhibit superior performance on downstream tasks compared to those pretrained on simple rules. The inclusion of both easy and hard tasks allows us to observe different performance trends and better understand the relationship between pretraining complexity, task difficulty, and emergent intelligence. The chess move prediction task, in particular, serves as an excellent system to test reasoning due to its inherent complexity and requirement for strategic thinking.
@@ 4.1 Downstream Task: Reasoning
Easy: In the easy task, we only use $3\times 3$ squares that are fixed in position and orientation. The only possible transformation is a color change, which we perform in a predetermined order.
Hard: The hard task involves more complex patterns where all transformations are applied simultaneously. We used four distinct base shapes, each represented by a $5\times 5$ matrix. This results in complex sequences where shapes change color, orientation, and position over time. The combination of these transformations requires the model to reason over multiple simultaneous changes to accurately predict the next pattern in the sequence.
For the easy reasoning task, models were post-trained for 1,000 epochs with a learning rate of $1\times 10^{-4}$. Due to the increased complexity and difficulty of the hard reasoning task, we post-train for 10,000 epochs with a learning rate of $1\times 10^{-5}$. We use early stopping based on validation loss to prevent overfitting.
@@ 4.2 Downstream Task: Chess Move Prediction
We added an embedding layer to convert the SAN tokens into vector representations, which were then processed using the frozen ECA-pretrained model. A linear output layer was used to transform the outputs to the vocabulary size corresponding to the SAN tokens. These input and output layers were trained while the rest of the model was frozen. The model was trained using cross-entropy loss, the Adam optimizer (Kingma, 2014), and a learning rate scheduler with warm-up. Early stopping was again used.
@@ 4.3 Hardware and Software
Figure 2: Relationship between downstream task performance and data complexity. (a) Eight representative ECA rules, two from each of Wolfram’s four complexity classes. Performance of models trained on these rules is highlighted in the top row of (b). (b) Top row: Model performance in relation to the Lempel-Ziv complexity of data generated by each rule. The left and center panels show efficiency (1 divided by number of epochs to reach 80% validation accuracy) for the easy and hard reasoning tasks, respectively. The right panel shows move prediction accuracy for the chess task. The rules depicted on the left are highlighted in the plot with triangles and annotated with the rule number. The correlation coefficient is shown in the top-left corner of each plot. An asterisk next to the value indicates a significant relationship ($p<0.05$). Bottom row: Downstream task performance based on Wolfram classification of each rule. Models trained on Class III and Class IV (chaotic and complex) rules perform better than models trained on uniform and simple rules. Baseline results for a randomly initialized transformer model are shown with a dashed black line on all plots.
@@ 5 Results
In this section, we present our results exploring the relationship between system complexity and emergent intelligence in LLMs. The following sections detail our analyses of task performance and attention patterns across models trained on data with varying complexities.
@@ 5.1 Relationship between Intelligence and Complexity
Figure 2 presents the model performance across three downstream tasks (easy reasoning, hard reasoning, and chess move prediction) as a function of the complexity of the ECA rules the models were pretrained on. The top row highlights the relationship between performance and the Lempel-Ziv complexity data, while the bottom row categorizes the performance by Wolfram’s complexity classes. For clarity, two representative rules from each complexity class are displayed on the left, with their corresponding performance annotated in the top plots.
For the reasoning tasks, models generally achieve near-perfect accuracy when trained for a sufficient number of epochs. Therefore, instead of reporting absolute accuracy, we focus on model efficiency, defined as the inverse of the number of epochs required to reach 80% accuracy. The chess task is sufficiently difficult that models do not achieve perfect performance, and so we report the final accuracy. As data complexity increases, we observe a clear positive correlation in all tasks, with more complex data leading to greater efficiency. This correlation is significant for each of the tasks. The p-values, along with the distance correlation coefficient, are given in Appendix D.
In terms of Wolfram’s classification, rules from Classes I and II (uniform and periodic) show lower average efficiency in the reasoning tasks compared to those from Classes III and IV (chaotic and complex). Class IV rules especially outperform the other classes on the chess move prediction task. This pattern suggests that models trained on more complex data tend to perform better on harder downstream tasks. Results with respect to other complexity measures are shown in Appendix E.
We observe that models trained on certain Class III (Chaotic) rules, such as Rules 105, 146, and 150, have poorer performance on the hard reasoning and chess move prediction tasks. This behavior is expected due to chaotic systems lacking the structured patterns necessary for effective learning. In other words, they may be too random to predict, leading to weaker downstream performance. Though we provide the correlation coefficient to quantify the relationship, the key takeaway is the qualitative insight: performance peaks at intermediate complexity and deteriorates with excessive complexity. These results highlight the existence of a “sweet spot” of complexity conducive to intelligence, where the system is still predictable yet hard to predict.
@@ 5.2 Models Learn Complex Solutions For Simple Rules
Figure 3: Attention scores for the final 10 states prior to the target state, showing that models trained on more complex data rely more heavily on past states for prediction. Left: Visualization of the last 10 states and the target state for representative rules from each of Wolfram’s complexity classes. Center: Attention scores for each of the last 10 states, highlighting that models trained on chaotic and complex (Class III and Class IV) rules focus more on recent states, while models trained on uniform rules exhibit consistently low attention. Periodic rules demonstrate a repeating attention pattern, suggesting that the model is learning to attend to earlier cycles of the same state rather than general state history. Right: Average attention across the final 10 states for all rules, plotted against Lempel-Ziv complexity. A strong positive correlation ($r=0.66$) indicates that models trained on higher complexity data attend more highly to historical states during prediction.
@@
Our findings reveal that models trained on rules that produce more complex dynamics tend to allocate higher attention to the last ten states, with a strong positive correlation ($r=0.66$) between data complexity and average attention. This suggests that models trained on complex data integrate information from past states to make their predictions. In contrast, models trained on simpler rules display lower attention across the last ten states. For example, Rule 168 (uniform) shows consistently low attention, indicating that previous states are not being utilized in the prediction process. Rule 179 (periodic) demonstrates a recurring pattern in its attention scores, where the model focuses on every other state. This behavior is explained by the nature of Rule 179, which produces an alternating pattern that repeats every two time steps (Figure 3). Thus, the model appears to be learning only this alternating cycle, rather than general state history. The simpler attention structure suggests that these models learn a trivial solution, which is in accordance with their poorer downstream performance.
@@
We emphasize that for both rules that generate simple data and those that generate complex data, a trivial solution exists: learning the instantaneous 8-bit rule and applying it to only the current state. Such a model would have no attention on previous states, as it only needs the current state to make its prediction. The fact that the complex models are attending to previous states indicate that they are learning a more complex solution to this simple problem, and we conjecture that this complexity is what makes the model “intelligent” and capable of repurposing learned reasoning to downstream tasks.
@@
Figure 4: Comparison of model performance on short-term (1-step) and long-term (5-step) prediction tasks for ECA rules. Points are colored by Lempel-Ziv complexity, with the dashed line indicating equal performance. Points below the line show better short-term performance.
@@
We had initially expected that predicting one step in the future would be too easy, and would result in every model learning the trivial (8-bit rule) solution. As such, we trained models to predict 5 steps ahead. Surprisingly, we found that models predicting only the next step not only learned non-trivial solutions, but even outperformed models predicting five steps ahead (Figure 4). This result suggests that even when predicting the immediate next state, the models are learning nontrivial solutions (provided that the underlying data is complex), capable of capturing complex patterns beyond the trivial rule-based approach.
@@
In this section, we explored the relationship between system complexity and emergent intelligence. Our results show that downstream model performance improves with pretraining on more complex data, but can deteriorate with excessive complexity or chaotic behavior, signifying a sweet spot of complexity conducive to intelligence. Attention patterns further reveal that models trained on complex data integrate historical information into their predictions, suggesting that they are learning more sophisticated solutions to relatively simple problems. We hypothesize that this complexity in the learned representations is a key factor enabling models to generalize and perform well on downstream tasks.